OpenStudy (anonymous):

Write each improper fraction as a mixed number and each mixed number as an improper fraction. 1.29/6 2.4 1/9 3.82/5 4.2 5/6

ganeshie8 (ganeshie8):

\(\frac{29}{6} = \frac{24+5}{6} \\ = \frac{24}{6} + \frac{5}{6} \\ = 4 + \frac{5}{6} \\ = 4\frac{5}{6}\)

ganeshie8 (ganeshie8):

see if that makes sense. for the 2nd problem, converting to improper fraction is pretty easy compared to whats shown above for converting to mixed fraction

OpenStudy (anonymous):

...........

ganeshie8 (ganeshie8):

hmm

OpenStudy (anonymous):

i need hellp!!!

ganeshie8 (ganeshie8):

im not good at understanding dot language, sorry.

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

if anyone can figure this out ur my best firend!